# PlaceFM: A Training-free Geospatial Foundation Model of Places using Large-Scale Point of Interest Data

### v2

@@ Abstract.
With the rapid growth and continual updates of geospatial data from diverse sources, geospatial foundation model pre-training for urban representation learning has emerged as a key research direction for advancing data-driven urban planning. Spatial structure is fundamental to effective geospatial intelligence systems; however, existing foundation models often lack the flexibility to reason about places, context-rich regions spanning multiple spatial granularities that may consist of many spatially and semantically related points of interest. To address this gap, we propose PlaceFM, a geospatial foundation model that captures place representations through a training-free, clustering-based approach. PlaceFM summarizes the entire point of interest graph constructed from U.S. Foursquare data, producing general-purpose region embeddings while automatically identifying places of interest. These embeddings can be directly integrated into geolocation data pipelines to support a variety of urban downstream tasks. Without the need for costly pre-training, PlaceFM provides a scalable and efficient solution for multi-granular geospatial analysis. Extensive experiments on two real-world prediction tasks, ZIP code–level population density and housing prices, demonstrate that PlaceFM not only outperforms most state-of-the-art graph-based geospatial foundation models but also achieves up to a 100× speedup in generating region-level representations on large-scale POI graphs. The implementation is available at https://github.com/mohammadhashemii/PlaceFM.
Geospatial Foundation Model, Graph-based Urban Representation Learning, General Purpose Encoders
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; August 31st-September 4th, 2026; Boston, MAisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Geographic information systemsccs: Information systems Location based servicesccs: Computing methodologies Neural networksccs: Computing methodologies Learning latent representations
1. Introduction
The analysis of geospatial large datasets unlocks a wide range of applications that support diverse social functions across today’s world. In particular, recent research has harnessed the power of geospatial data to predict disease outbreaks (Kohn et al., 2023), detect behavioral patterns and anomalies (Amiri et al., 2024; Zhang et al., 2024a, b), monitor and forecast traffic, and inform smarter urban planning (Mokbel et al., 2024, 2022). Numerous machine learning methods have been developed across diverse data modalities, such as satellite imagery, search queries, economic indicators, and influenza trends, to learn informative representations of geospatial data (Du et al., 2020; Rolf et al., 2021; Ommi and Hashemi, 2024). In addition to these task-specific models, a range of prior approaches have also aimed to develop general-purpose geographic encoders that can be applied across multiple downstream geospatial tasks (Klemmer et al., 2025; Vivanco Cepeda et al., 2023). Despite their successes, existing machine learning approaches have often been limited by their task-specific nature, requiring carefully curated datasets and model architectures that generalize poorly across domains (Bommasani et al., 2021). This limitation has motivated the emergence of foundation models, large-scale pre-trained models that learn transferable representations from massive and diverse datasets. Foundation models first rose to prominence in natural language processing with the advent of large language models, and later transformed computer vision and multimodal models (Devlin et al., 2019; Brown et al., 2020; Radford et al., 2021). Their ability to generalize across tasks, modalities, and domains has led to widespread adoption in fields ranging from healthcare and biology to law and finance (Yang et al., 2022; Thirunavukarasu et al., 2023). Inspired by these advances, researchers have begun to explore geospatial foundation models and general-purpose encoders that aim to overcome the limitations of task-specific geospatial encoders (Liang et al., 2025).
Recently, several geospatial foundation models have been developed that focus on geolocation representation learning, aiming to generate general-purpose embeddings for geographic entities, such as urban regions (Liang et al., 2025; Tempelmeier et al., 2021; Li et al., 2022; Wu et al., 2023). Urban regions, as spatially distributed neighborhoods with relatively homogeneous physical and socioeconomic characteristics (Yuan et al., 2012), serve as a natural analytical unit for tasks such as region function recognition (Zhang et al., 2017, 2018), and population estimation (Cheng et al., 2022). By representing urban regions through embeddings learned from large-scale geospatial data, such as Points of Interest (POI) data or satellite images, these foundation models enable scalable and transferable analysis across multiple urban studies, bridging the gap between fine-grained location embeddings and aggregated urban patterns. In this context, it is reasonable to expect that learning meaningful representations for urban regions, capable of capturing and disentangling the underlying factors encoded in raw geolocation data, can fundamentally enhance a wide range of urban downstream tasks analyses (Mai et al., 2020).
@@ 1. Introduction
To derive meaningful representations of urban regions from raw POI data, a common approach is to construct a POI graph in which nodes correspond to individual POIs and edges are defined using a geographical proximity function. General-purpose encoders are then trained on this graph to generate informative embeddings that capture the characteristics of urban regions(Jin et al., 2023; Yan et al., 2017; Agarwal et al., 2024; Huang et al., 2023). For instance, PDFM (Agarwal et al., 2024) pretrains on diverse geospatial data sources such as maps, activity levels, search trends, weather, and air quality by constructing a heterogeneous graph where nodes represent counties and postal codes and edges capture spatial proximity. A graph neural network (GNN) is then applied to learn embeddings for these geographic units. Similarly, HGI (Huang et al., 2023) aggregates POI data to the regional level with a multi-head attention mechanism, applies graph convolution to model spatial dependencies, and hierarchically combines these regional embeddings into city-level representations. Despite their potential, current geospatial foundation models for urban representation learning suffer from three key limitations that hinder their effectiveness in real-world applications:
@@
(a) The Missing Sense of Place notion: While geospatial foundation models have advanced urban representation learning and demonstrated effectiveness across a variety of downstream tasks, current methods still struggle to capture the notion of places, regions shaped by human meaning and behavior that consist of spatially and semantically related POIs (Hashemi and Zufle, 2025). In other words, most existing methods produce learned representations at a fixed geographic granularity, such as ZIP codes, cities, and assign a single embedding to represent an entire region. The black-box process of generating a single vector to represent an entire region overlooks the fact that within any given area, multiple semantically meaningful subregions may exist that are highly relevant to users interests but not necessarily aligned with administrative boundaries. For example, a “cat lovers” place could emerge organically around a cluster of a cat-themed café, a park frequented by stray cats, and nearby pet shops, yet such meaningful patterns are lost when restricted to rigid spatial units. Moreover, representations derived solely from individual POIs, while useful as fine-grained descriptors of the built environment, fail to capture the broader semantic context that gives a region its true meaning (Niu and Silva, 2021).
(b) High Pre-training Computational Cost: A key challenge of existing geospatial foundation models for urban representation learning lies in the computational cost of the pre-training stage, which requires learning from massive POI datasets. Most current methods pre-train their encoders on a limited number of POIs, often by selecting sample cities or regions, which simplifies training but does not reflect real-world scales (Huang et al., 2023). Scaling to a nationwide dataset, for example, over twenty million POIs in the United States from Foursquare [^1] https://docs.foursquare.com/data-products/, would impose substantial computational overhead, demanding significant training time and resources while still aiming to learn effective and generalizable representations.
@@
(c) Lack of Granularity Flexibility: Current geospatial foundation models are typically constrained to a single level of spatial granularity during both pre-training and inference time, restricting their flexibility in downstream applications. For example, PMT (Wu et al., 2024) encodes trajectories as sequences of United States Census Block Groups (CBGs), with all subsequent tasks, such as next-location prediction, restricted to that level. Similarly, PDFM (Agarwal et al., 2024) produces embeddings only for ZIP codes and counties in the United States, preventing inference at finer scales such as neighborhoods or CBGs. In practice, however, the notion of a place can be understood at multiple levels of granularity depending on the task and user perspective(Hashemi and Zufle, 2025) and may not align with political and administrative boundaries. For example, POIs related to a “cat lovers” place may span across multiple ZIP-code and CBG boundaries.
To address these three critical challenges, we propose PlaceFM, a training-free geospatial foundation model that is simple yet effective. PlaceFM is capable of (1) Multi-granular Region Representation Learning: generating meaningful representations for geographic entities across multiple levels of granularity, and (2) Place Identification: identifying places composed of spatially and semantically related POIs.
We begin by formally defining the notion of a place and outlining the motivation for identifying places within a geographical region composed of POIs. Subsequently, we introduce FSQ-19M, a large-scale POI graph dataset comprising approximately 19 million POIs constructed from the Foursquare tabular dataset, spanning all 48 contiguous U.S. states. Building on this resource, we propose a semantic-spatial-aware clustering-based approach for generating region embeddings while simultaneously identifying meaningful places within each region. Notably, PlaceFM enhances traceability by establishing an explicit correspondence between identified places and their original POIs, thereby providing deeper insights into how these places encapsulate information from the underlying POI graph. In summary, our contributions are:
@@
  • For the first time, we formalize the notion of a place in geospatial POI graph data and propose the first framework that automatically identifies places within a regional graph. The framework partitions POI representations into clusters, where the resulting centroids serve as place node embeddings.
@@
In Section 2, we review existing geospatial foundation models and general-purpose encoders for urban representation learning, highlighting their key challenges and limitations. Section 3 then formalizes the problem setting, introduces the definition of a place, and outlines the scope of our study. The proposed methodology, including data integration strategies, model design, and system architecture, is detailed in Section 4. Section 5 presents the experimental setup, evaluation metrics, and performance results. Finally, Section 6 concludes the paper and outlines future research directions.
2. Related Works
2.1. Geospatial Foundation Models
Pre-trained geospatial foundation models are large-scale neural networks trained on diverse geospatial datasets, enabling them to capture transferable representations with broad utility across regions and tasks (Liang et al., 2025). Unlike traditional machine learning models that often fail to generalize beyond localized settings, these models can be effectively fine-tuned with limited task-specific data, offering scalability, adaptability, and efficiency for a wide range of spatial analysis applications. To address the issues in traditional machine learning approaches for geospatial data analysis, recent efforts have introduced geospatial foundation models designed for geolocation representation learning, with the goal of producing general-purpose embeddings for geographic entities (Liang et al., 2025; Agarwal et al., 2024; Balsebre et al., 2024; Tempelmeier et al., 2021; Li et al., 2022; Yan et al., 2024; Xiao et al., 2024; Wu et al., 2023). These embeddings serve as a backbone for a wide range of geospatial downstream tasks, facilitating more scalable and effective model development. For instance, GeoVectors (Tempelmeier et al., 2021) and SpaBERT (Li et al., 2022) leverage OpenStreetMap data [^2] https://www.openstreetmap.org/ to learn location embeddings, while G2PTL (Wu et al., 2023) is pre-trained on large-scale logistics delivery data.
Large-scale POI datasets have gained attention for pretraining geospatial foundation models(Jin et al., 2023). Being inherently multimodal, they combine signals such as satellite imagery, textual reviews, categorical attributes, and mobility patterns, enabling representations that capture both spatial and semantic properties. POI data can also be modeled as graph-structured data, where nodes represent locations and edges capture spatial or functional relationships (Chen et al., 2022). One of the closest related works, PDFM (Agarwal et al., 2024), frames geospatial pretraining as a graph learning problem. It constructs a heterogeneous graph where nodes represent counties and postal codes enriched with diverse signals such as maps, activity levels, search trends, weather, and air quality, while edges encode spatial proximity. A GNN is then employed to learn embeddings that capture complex geographic relationships. Despite the effectiveness of its embeddings in downstream tasks, PDFM is limited to generating general-purpose representations only at the ZIP code and county levels, restricting its ability to capture places at finer granularities such as neighborhoods or CBGs.
Figure 1. The spatial distribution of POIs based on a uniform random sample of one million entries from the FSQ-19M dataset, covering the 48 contiguous U.S. states.
2.2. Graph-based Urban Representation Learning
Urban representation learning aims to encode geographical entities, such as POIs and neighborhoods, into meaningful vector representations that capture both the functional and structural characteristics of cities. Such representations are crucial for understanding urban dynamics and supporting downstream tasks, including land use classification, location recommendation, and urban planning (Zhang et al., 2024c; Liang et al., 2025). Early approaches focused on learning embeddings from POIs to summarize the functional properties of geographical regions. Inspired by Word2Vec(Church, 2017), in (Yao et al., 2017a), embeddings for POI categories are learned based on co-occurrence patterns, while Place2Vec (Yan et al., 2017), improved this by leveraging K-nearest-neighbors (KNN) to capture contextual similarities between POIs. Subsequent works extended these methods to generate region-level representations using POI category embeddings(Zhai et al., 2019; Niu and Silva, 2021). However, these approaches often ignored the uniqueness of individual POIs: for instance, two restaurants in different spatial contexts would receive identical category embeddings, overlooking subtle yet meaningful differences. Recent studies addressed this limitation by incorporating the spatial context of each POI using GNN and message passing-based algorithms, defining region representation as a supervised graph classification problem (Xu et al., 2022). Modern urban representation learning methods have increasingly utilized graph-based approaches to capture relationships between regions. HGI (Huang et al., 2023) exemplifies this trend by aggregating POIs to the regional level using multi-head attention to model their diverse influences, applying graph convolution to encode similarities between adjacent regions, and further aggregating these regional representations into a city-level embedding. Such graph-centric models allow for more flexible and context-aware representations compared to methods restricted to fixed geographic units or category-level embeddings; however, realizing this potential, for example, in HGI, whose encoder is based on an attention-based neural network, requires training on massive POI datasets, incurring substantial computational overhead.
3. Preliminaries & Definitions
In this section, we provide a detailed description of all the notations and definitions:
  • $\mathcal{P}$: Set of all Points of Interest (POIs), where each POI $p\in\mathcal{P}$ is represented as $(\text{lat}_{p},\text{lon}_{p},A_{p})$, with $A_{p}=[a_{p_{1}},a_{p_{2}},\dots,a_{p_{K}}]$ denoting the list of $K$ attributes of $p$.
  • $a_{p_{i}}$: The $i$-th attribute of a POI $p$. In our study, we only use the hierarchical category attribute, though other attributes could be incorporated. As an example, $\text{category}_{p}$ as the hierarchical category attribute can be: ”Dining and Drinking $\rightarrow$ Middle Eastern Restaurant $\rightarrow$ Persian Cafe”.
  • $G=(\mathcal{V},\mathcal{E})$: Graph of POIs, where $\mathcal{V}=\mathcal{P}$ is the set of nodes, $\mathcal{E}$ is the set of edges, and edges are defined based on a proximity function $d(p_{i},p_{j})$.
  • $G_{r}=(\mathcal{V}_{r},\mathcal{E}_{r})$: Urban region, represented as a subgraph of POIs, with $\mathcal{V}_{r}\subseteq\mathcal{V}$ and $\mathcal{E}_{r}\subseteq\mathcal{E}$.
3.1. Place Definition
@@
Developing a geospatial foundation model requires moving beyond point-based map representations toward a structured concept of places (Hashemi and Zufle, 2025). Here, we define a place as a semantically meaningful spatial unit that may align with or encompass multiple geographic entities, including POIs, postcodes, neighborhoods, or administrative regions.
Definition 3.1 (Place).
@@
$\mathcal{E}=P\cup\mathcal{P}$ denotes the universe of geographic entities, including primitive elements $\mathcal{P}$ (i.e., POIs) and higher-level places $P$ that are both semantically and spatially similar. This recursive definition allows us to capture the hierarchical nature of places, where smaller places can be nested within larger ones. The similarity between elements can be computed using any appropriate function, such as Euclidean distance or cosine similarity. The location of each entity $e_{i}$ is described by their coordinates $(\text{lat}_{e_{i}},\text{lon}_{e_{i}})$, where
[…]
3.2. Problem Statement
@@
3.2.1. Region-level representation learning
The first objective is to model urban environments across multiple spatial scales by compressing large-scale POI-level data into compact and informative region-level embeddings. Formally, for each region $G_{r_{i}}$ we seek to learn a $d$-dimensional feature vector $z_{r_{i}}\in\mathbb{R}^{d}$ such that
$z_{r_{i}}=f(G_{r_{i}}),$
where $f:\mathcal{G}\to\mathbb{R}^{d}$ is a representation function mapping graphs of POIs into a low-dimensional latent space. The generated embeddings $\{z_{r_{i}}\}_{i=1}^{N}$ should be both generalizable and effective across a wide range of downstream urban analytics tasks.
3.2.2. Place Discovery within Regions.
The second objective is to automatically identify higher-level places within each urban region $G_{r_{i}}$. A place $P\subseteq G_{r_{i}}$ as in Definition 3.1 is a subgraph consisting of POIs ro higher-level places that are both semantically and spatially similar. For each place $P$, we aim to compute an embedding
$z_{P}=f(P),$
such that $z_{P}$ captures the collective semantic and spatial characteristics of its constituent POIs. This hierarchical formulation ensures traceability, since we can explicitly associate each POI with its corresponding place, and compare similarity across places both within and across regions. Such capability has practical applications in place discovery (e.g., identifying emerging functional areas) and place recommendation (e.g., detecting multiple “cat-lovers” neighborhoods in a city).
Overall, the goal is to learn a representation function $f$ that supports hierarchical and multi-scale modeling of geographic space. This enables flexible analysis at different levels of granularity, from individual POIs to regions, while preserving the semantic interpretability and spatial coherence of discovered places.
4. Methodology
In this section, we provide a detailed overview of the data preparation process, the construction of the POI-level graph, and the explanation of the foundation model proposed in this paper. The architecture overview of our proposed method, PlaceFM, is depicted in Figure 2.
Figure 2. Pipeline of the proposed geospatial foundation model, PlaceFM. First, it builds POI-level graphs for each state, followed by category feature encoding and feature propagation to obtain neighborhood-aware POI embeddings. Places are then identified via training-free clustering at a chosen granularity, and an aggregator function produces the final region-level embedding.
4.1. FSQ-19M POI Dataset Preparation

### v3

# Training-Free Geospatial Place Representation Learning from Large-Scale Point-of-Interest Graph Data

@@ Abstract
Learning effective representations of urban environments requires capturing spatial structure beyond fixed administrative boundaries. Existing geospatial representation learning approaches typically aggregate Points of Interest(POI) into pre-defined administrative regions such as census units or ZIP code areas, assigning a single embedding to each region. However, POIs often form semantically meaningful groups that extend across, within, or beyond these boundaries, defining places that better reflect human activity and urban function. To address this limitation, we propose PlaceRep, a training-free geospatial representation learning method that constructs place-level representations by clustering spatially and semantically related POIs. PlaceRep summarizes large-scale POI graphs from U.S. Foursquare data to produce general-purpose urban region embeddings while automatically identifying places across multiple spatial scales. By eliminating model pre-training, PlaceRep provides a scalable and efficient solution for multi-granular geospatial analysis. Experiments using the tasks of population density estimation and housing price prediction as downstream tasks show that PlaceRep outperforms most state-of-the-art graph-based geospatial representation learning methods and achieves up to a $100\times$ speedup in generating region-level representations on large-scale POI graphs. The implementation of PlaceRep is available at https://github.com/mohammadhashemii/PlaceRep.
1 Introduction
The analysis of large-scale geospatial data enables a wide range of applications that support diverse social and urban functions. A broad range of machine learning methods have been developed to learn representations from diverse geospatial data modalities, including satellite imagery, search queries, economic indicators, and mobility-related signals (Du et al., 2020; Rolf et al., 2021; Ommi and Hashemi, 2024). Despite their successes, existing machine learning approaches have often been limited by their task-specific nature, requiring carefully curated datasets and model architectures that generalize poorly across domains (Bommasani et al., 2021). A key step toward addressing this limitation is to identify appropriate spatial regions over which transferable geospatial representations can be learned and reused across tasks. These geospatial regions, defined as spatially distributed neighborhoods with relatively homogeneous physical and socioeconomic attributes (Yuan et al., 2012), form a natural analytical unit for various urban planning tasks (Zhang et al., 2017, 2018; Cheng et al., 2022). Learning region-level embeddings from large-scale geospatial data, including POI data and satellite images, enables scalable and transferable analysis across diverse urban studies and downstream tasks (Liang et al., 2025). These embeddings also serve as reusable building blocks for geospatial foundation models that require transferable representations across tasks and spatial contexts (Liang et al., 2025; Tempelmeier et al., 2021; Li et al., 2022; Wu et al., 2023).
@@ 1 Introduction
To derive meaningful representations of geospatial regions from raw POI data, a common approach is to construct a POI graph in which nodes correspond to individual POIs and edges are defined using a geographical proximity function. General-purpose encoders are then trained on this graph to generate informative embeddings that capture the characteristics of urban regions(Jin et al., 2023; Yan et al., 2017; Agarwal et al., 2024; Huang et al., 2023). Despite their potential, current geospatial representation learning models suffer from three key limitations that hinder their effectiveness in real-world applications:
@@
(a) The Missing Sense of Place notion: Existing methods struggle to capture the notion of places, regions shaped by human meaning and behavior that consist of spatially and semantically related POIs (Hashemi and Zufle, 2025). Most approaches learn representations at a fixed geographic granularity, assigning a single embedding to an entire region and treating it as a black box. This overlooks the fact that multiple semantically meaningful subregions may exist within the same area and may not align with administrative boundaries. For example, a “cat lovers” place may emerge around a cat-themed café, nearby parks frequented by stray cats, and surrounding pet shops, yet such patterns are lost under rigid spatial units. Moreover, representations based solely on individual POIs, while useful at fine spatial scales, fail to capture the broader semantic context that gives a region its meaning (Niu and Silva, 2021).
(b) High Pre-training Computational Cost: A major limitation of existing geospatial region representation methods is the high computational cost of pretraining on large POI datasets. To reduce complexity, most approaches train on sampled cities or regions, which limits scalability and realism (Huang et al., 2023). Extending these methods to nationwide datasets with tens of millions of POIs would incur substantial training time and resource demands.
@@
(c) Lack of Granularity Flexibility: Current geospatial region representation learning models are typically constrained to a single level of spatial granularity during both pre-training and inference time, restricting their flexibility in downstream applications. For example, PMT (Wu et al., 2024) encodes trajectories as sequences of United States Census Block Groups (CBGs), with all subsequent tasks, such as next-location prediction, restricted to that level. Similarly, PDFM (Agarwal et al., 2024) produces embeddings only for ZIP codes and counties in the United States, preventing inference at finer scales such as neighborhoods or CBGs.
To address these three critical challenges, we propose PlaceRep capable of (1) Multi-granular Region Representation Learning: generating meaningful representations for geographic entities across multiple levels of granularity, and (2) Place Identification: identifying places composed of spatially and semantically related POIs. We first define the concept of a place and motivate its identification within POI-based geographic regions. We then introduce FSQ-19M, a large-scale POI graph dataset with approximately 19 million POIs spanning the 48 contiguous U.S. states. Building on this dataset, we propose a semantic–spatial clustering approach to jointly learn region embeddings and identify meaningful places, while preserving traceability through explicit mappings between places and their constituent POIs. Our contributions are summarized as follows:
@@
  • We formalize the notion of a place in geospatial POI graph data and propose the first framework that automatically identifies places within a regional graph. The framework partitions POI representations into clusters, where centroids serve as place node embeddings.
@@
2 Related Works
Geospatial region representation learning seeks to encode geographic entities such as POIs, neighborhoods, and regions into vector embeddings that capture both functional semantics and spatial structure, enabling downstream tasks including land-use classification, recommendation, and urban analysis (Zhang et al., 2024; Liang et al., 2025). Early approaches learned POI or category embeddings based on co-occurrence patterns or spatial proximity, inspired by Word2Vec (Yao et al., 2017a; Yan et al., 2017), and were later extended to derive region-level representations by aggregating POI category embeddings (Zhai et al., 2019; Niu and Silva, 2021). However, these methods often overlooked the unique spatial context of individual POIs.
Recent work addresses this limitation by modeling POIs and regions as graph-structured data and leveraging GNNs or message-passing mechanisms to capture contextual and relational information (Xu et al., 2022; Huang et al., 2023). In parallel, large-scale multimodal POI datasets have enabled pretraining geospatial foundation models that integrate signals such as satellite imagery, textual attributes, and mobility patterns (Jin et al., 2023). Among the closest related efforts, PDFM (Agarwal et al., 2024) formulates geospatial pretraining as heterogeneous graph learning over regions enriched with diverse environmental and activity signals, learning general-purpose embeddings via GNNs. While effective for downstream tasks, PDFM operates at coarse spatial units such as ZIP codes and counties, limiting its ability to represent finer-grained urban regions such as neighborhoods.
3 Preliminaries & Definitions
In this section, we present the formal definitions and problem formulation. Notation used throughout the paper is summarized in the Appendix B.
3.1 Place Definition
@@
Developing a geospatial representation learning method requires moving beyond point-based map representations toward a structured concept of places (Hashemi and Zufle, 2025). Here, we define a place as a semantically meaningful spatial unit that may align with or encompass multiple geographic entities, including POIs, postcodes, neighborhoods, or administrative regions.
@@
$\mathcal{E}=P\cup\mathcal{P}$ denotes the universe of geographic entities, including primitive elements $\mathcal{P}$ (i.e., POIs) and higher-level places $P$ that are both semantically and spatially similar. This recursive definition allows us to capture the hierarchical nature of places, where smaller places can be nested within larger ones. The location of each entity $e_{i}$ is described by their coordinates $\text{loc}_{e_{i}}$:
$\text{loc}_{e_{i}}=\begin{cases}(\text{lat}_{p_{i}},\text{lon}_{p_{i}}),&\text{if $e_{i}\in\mathcal{P}$,}\\ \text{centroid}\big(\{(\text{lat}_{p_{j}},\text{lon}_{p_{j}})\mid p_{j}\in e_{i}\}\big),&\text{else.}\end{cases}$ (1)
3.2 Problem Statement
@@
3.2.1 Region-level representation learning
The first objective is to model urban environments across multiple spatial scales by compressing large-scale POI-level data into compact and informative region-level embeddings. Formally, for each region $G_{r_{i}}$ we seek to learn a $d$-dimensional feature vector $z_{r_{i}}\in\mathbb{R}^{d}$ such that $z_{r_{i}}=f(G_{r_{i}})$, where $f:\mathcal{G}\to\mathbb{R}^{d}$ is a representation function mapping graphs of POIs into a low-dimensional latent space. The generated embeddings $\{z_{r_{i}}\}_{i=1}^{N}$ should be both generalizable and effective across a wide range of downstream urban analytics tasks.
3.2.2 Place Discovery within Regions
The second objective is to automatically identify higher-level places within each urban region $G_{r_{i}}$. A place $P\subseteq G_{r_{i}}$ as defined in Section 3.1 is a subgraph consisting of POIs ro higher-level places that are both semantically and spatially similar. For each place $P$, we compute an embedding $z_{P}=f(P)$, such that $z_{P}$ captures the collective semantic and spatial characteristics of its respective POIs. This hierarchical formulation ensures traceability, since we can explicitly associate each POI with its corresponding place, and compare similarity across places both within and across regions. Such capability has applications in place discovery (e.g., identifying emerging functional areas) and place recommendation (e.g., detecting “cat-lovers” neighborhoods in a city).
4 Methodology
In this section, we provide a detailed overview of the data preparation process, the construction of the POI-level graph, and the explanation of the gepspatial representation learning model proposed in this paper. The architecture overview of our proposed method is depicted in Figure 1.
Figure 1: PlaceRep’s architecture overview. First, it builds POI-level graphs for each state, followed by category feature encoding and feature propagation to obtain neighborhood-aware POI embeddings. Places are then identified via training-free clustering at a chosen granularity, and an aggregator function produces the final region-level embedding.
4.1 FSQ-19M POI Dataset Preparation